40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 846     Entry time: Mon Aug 18 11:50:29 2008
Author: Yoichi 
Type: Update 
Category: SUS 
Subject: In vacuum free swinging results 
The first attachment is the results of the free swinging spectra measurement performed in vacuum this morning.
They are freely swinging, but the suspensions in the BS chamber got even more extra peaks.
Especially, the SRM spectrum looks like a forest.
If those extra peaks are inter-modulations of the primary suspension modes, the heights of them should be
enhanced (compared to the in-the-air case) by the increased quality factors of the primary modes (due to the less air friction).
This might explain the observed increase in the extra peaks.

While doing the free swinging, we had two big spikes in the OSEM signals of the ETMs and only in ETMs.
Those spikes screwed up the spectra of the ETMs. So the ETM spectra were calculated using the time series
after the spikes.
The second attachment shows one of those spikes. It looks like a computer glitch.
Attachment 1: 2008-8-18.pdf  1.546 MB  | Hide | Hide all
2008-8-18.pdf 2008-8-18.pdf 2008-8-18.pdf 2008-8-18.pdf 2008-8-18.pdf 2008-8-18.pdf 2008-8-18.pdf 2008-8-18.pdf
Attachment 2: spike.pdf  134 kB  | Hide | Hide all
spike.pdf
ELOG V3.1.3-